On June 1, 2019 Adelphi Corporation issued $380,000 of 6%, 5-year bonds.  The bonds which were issued at 95, pay interest on January 1 and June 1. Use this information to calculate the amount of bond discount or premium that is amortized with each interest payment. If this is discount amortization enter as a positive number. If this is premium amortization enter as a negative number.
Answer #1

Discount on bonds payable = 380000*5/100 = 19000

Bond discount amortized with each interest payment = 19000/10 = 1900 per semiannual period